the perimeter of a shape is 275 feet. 4 of its 6 sides add up to 195 feet. The two remaining sides are equal. What is the length of each remaining side?
Answer by ankor@dixie-net.com(22740)

the perimeter of a shape is 275 feet. 4 of its 6 sides add up to 195 feet.
The two remaining sides are equal.
What is the length of each remaining side?
:
let s = the length of each remaining side
195 + 2s = 275
2s = 275 - 195
2s = 80
s = 80/2
s = 40 ft